OpenStudy (anonymous):

What is -6n = 11/12 ??

OpenStudy (anonymous):

How do you undo multiplication?

OpenStudy (idkwut):

\[-6n = 11/12\] Do you know how to undo multiplication?

OpenStudy (anonymous):

The opposite of multiplication is division. So divide by -6

OpenStudy (anonymous):

Hello :) for your question you have to isolate (separate something from another thing) the variable from the constants. Since you have -6n = 11/12, dividing each side by -6 is the same thing as multiplying both sides by 1/-6. So you get -11/72 = n

OpenStudy (anonymous):

If you would like this written down in mathematical talk rather than a paragraph let me or one of these other kind persons know. :)

OpenStudy (anonymous):